OpenStudy (anonymous):

9^-1/5. Write in Radical Form.

OpenStudy (anonymous):

Firstly, \[\huge x^{-\frac{a}{b}} \implies \sqrt[b]{x^{-a}}\] and: \[\huge x^{-a} = \frac{1}{x^a}\]

OpenStudy (anonymous):

So I have to find the radical first then, place it in a form without the negative exponent?

OpenStudy (anonymous):

Yep...

OpenStudy (anonymous):

Or you can leave it in negative form in the radical also, there is no harm..

OpenStudy (anonymous):

So whats the answer?

OpenStudy (anonymous):

You tell me first, I will check it for you..

OpenStudy (anonymous):

Is it \[\frac{ 1 }{ \sqrt[5]{9} }\]

OpenStudy (anonymous):

Oh sorry, YES IT IS RIGHT..

OpenStudy (anonymous):

Oh. Lol Thank you so much then, just looing for confirmation, This homework doesnt have an answer sheet,